# Equilibrium Cycle: A “Dynamic” Equilibrium

### v1

@@ Abstract
The Nash equilibrium (NE) is fundamental game-theoretic concept for characterizing stability in static strategic form games. However, at times, NE fails to capture outcomes in dynamic settings, where players’ actions evolve over time in response to one another. In such cases, game dynamics fail to converge to an NE, instead exhibiting cyclic or oscillatory patterns. To address this, we introduce the concept of an ‘equilibrium cycle’ (EC). Unlike NE, which defines a fixed point of mutual best responses, an EC is a set-valued solution concept designed to capture the asymptotic or long-term behavior of dynamic interactions, even when a traditional best response does not exist. The EC identifies a minimal rectangular set of action profiles that collectively capture oscillatory game dynamics, effectively generalizing the notion of stability beyond static equilibria. An EC satisfies three important properties: stability against external deviations (ensuring robustness), unrest with respect to internal deviations (driving oscillation), and minimality (defining the solution’s tightness). This set-valued outcome generalizes the minimal curb set to discontinuous games, where best responses may not exist. In finite games, the EC also relates to sink strongly connected components (SCCs) of the best response graph.
@@
Cyclic Behaviour,
Economic Games,
Nash Equilibrium,
@@ 1 Introduction
However, there are well known issues with applying the Nash equilibrium to understand game dynamics. By game dynamics, we mean settings where players change their actions over time, seeking to improve their own payoff, in response to the actions of other players. Such game dynamics, which includes the special case of best response dynamics, do not always converge to a Nash equilibrium (Demichelis et al., 2003; Benaïm et al., 2012). From a dynamical systems standpoint, this is not surprising. After all, NEs are simply stationary points (a.k.a., equilibria) corresponding to a broad class of game dynamics, and dynamical systems do not in general converge to a stationary point. Indeed, it is not unusual for game dynamics to result in limit cycles (i.e., the action profile oscillates asymptotically), or even more chaotic behavior in higher dimensions; see Benaïm and Hirsch (1999); Ficici et al. (2005); Benaïm et al. (2012); Papadimitriou and Piliouras (2019). In such situations, the Nash equilibrium as a solution concept does not meaningfully capture the ‘outcome’ of the (dynamic) strategic interaction between the players.
In this paper, we introduce a novel solution concept, which we call the equilibrium cycle (EC), that seeks to capture the outcome of oscillatory game dynamics. Specifically, we restrict attention to pure (i.e., non-randomised) player strategies. The EC seeks to capture the limit set associated with a broad class of such (oscillatory) game dynamics. Crucially, the definition of the EC does not require the existence of best responses, and is therefore also applicable to discontinuous games, which arise naturally in various contexts (Dasgupta and Maskin, 1986a, b; Reny, 1999).
@@
It is easy to observe that this game does not have a pure NE (see Lotker et al. (2008)). Let us now consider natural ‘near-best response’ dynamics on this game. Note that best responses per se do not exist in this game owing to discontinuities in the payoff function. By near-best response, we mean an action that yields nearly-maximal utility for a player in response to the opponent’s action.
@@ 2 Equilibrium Cycle: Definition and examples
Observe that $0<b\leq 0.5,$ since $F$ is closed and ${\mathcal{F}}\cap{\mathcal{D}}\neq\emptyset$. We now consider the following sub-cases:
@@ 1. $(i)$
    $\{x\in[0,b)\colon(x,x)\in{\mathcal{F}}\}=\emptyset:$ Here, by the construction of $b,$ it follows that $(0,0)\notin{\mathcal{F}}$ and $(0.5,0.5)\in{\mathcal{F}}.$ This contradicts the first condition for an EC, since the unique best response of any player $i$ to the opponent’s action $a_{-i}=0.5$ is $0\notin{\mathcal{F}}_{i}.$
@@ 2. $(ii)$
    By the construction of $b$ and $c$, it follows that $c<b,$ $(c,c)$ lies in ${\mathcal{F}},$ and $(x,x)\notin{\mathcal{F}}$ for all $x\in(c,b).$ Now, for any player $i,$ say $a_{-i}=c.$ Then for $a_{i}\in(c,b),$
@@
It is instructive to note that in the classical Bertrand duopoly, $O_{c}=0,$ resulting in the unique Nash equilibrium $(c,c).$ It is the introduction of a positive operational cost (that is not so large that it prohibits firms from operating at all, as ensured by our assumption) that results in the EC characterized in Lemma 3. Finally, we note that the linear demand assumption is not essential to the emergence of the EC; an analogous EC can be established under any smooth decreasing demand function (so long as a positive utility is possible). The proof of Lemma 3 is analogous to that of Lemma 2, and is therefore omitted.
The Bertrand duopoly posits an extreme ‘all or nothing’ bifurcation of market demand between the firms. The following example, which generalizes a pricing game between competing ride-hailing platforms in Walunj et al. (2023), considers a ‘softer’ payoff discontinuity, a more realistic setting.
@@
The game in Example 3 also admits an EC, as shown in the following lemma (proof is once again omitted, given its similarity to the proof of Lemma 2).
@@ 5 Concluding remarks
In classical game theory, the Nash equilibrium (Nash Jr, 1950) is a foundational concept, identifying strategy profiles where no player can unilaterally deviate to improve their payoff, making it ideal to describe static one-shot interactions, and also convergent game dynamics. Indeed, when the outcome of game dynamics is captured by a single point (strategy profile), it is typically a (pure) Nash equilibrium. However, when this is not the case, more generalized solution concepts are needed. For example, it might be meaningful to identify a set of points (strategy profiles) that become relevant in the long run, i.e., in the limit. In other words, we might anticipate that well-known game dynamics, such as best/better response dynamics, would converge to these sets over time, which can then be considered as the outcome of the game.
@@
One such set-valued outcome of game dynamics is the minimal curb set, introduced in Basu and Weibull (1991). These sets have become essential for studying adaptive processes in games, as many adjustment processes naturally settle within a minimal curb set (Voorneveld et al., 2005; Hurkens, 1995). However, this notion assumes the existence of best responses. In this paper, we define an alternative set-valued equilibrium notion, the equilibrium cycle, that does not assume the existence of best responses, and can meaningfully be applied even in discontinuous games. The equilibrium cycle is a minimal rectangular (more formally, Cartesian product) set of action profiles that is stable against external deviations, while also being unstable with respect to internal deviations. Interestingly, minimal curb sets do not possess the latter ‘unrest’ property; however, non-trivial minimal curb sets, which exclude pure Nash equilibria, do, and are equivalent to equilibrium cycles in best response games. Importantly, payoff discontinuities arise naturally in several economics applications (Dasgupta and Maskin, 1986a, b; Reny, 1999); these discontinuous games tend not to have best responses. In this regard, the equilibrium cycle represents a significant generalization over the curb set on non-BR games, as illustrated by several examples in this paper.
@@
At the other extreme, specializing to finite games, we find that the equilibrium cycle is related to strongly connected sink components (or sink SCCs) of the best response graph. However, an exact equivalence does not hold here, owing primarily to the implicit rectangularity of equilibrium cycles (the same rectangularity does not hold in general for sink SCCs). It is possible to define a ‘non-rectangular’ variation of the equilibrium cycle that is equivalent to sink SCCs of the best response graph; however, this variant is difficult to apply to discontinuous games, and also loses the equivalence to (non-trivial, minimal) curb sets in BR games. In particular, we find that both the rectangularity as well as the closed-ness of the EC are essential for application in discontinuous games, such as those describes in Examples 13.

### v2

@@ Abstract
In this paper, we introduce a novel equilibrium concept, called the equilibrium cycle, which seeks to capture the outcome of oscillatory game dynamics. Unlike the (pure) Nash equilibrium, which defines a fixed point of mutual best responses, an equilibrium cycle is a set-valued solution concept that can be demonstrated even in games where best responses do not exist (for example, in discontinuous games). The equilibrium cycle identifies a Cartesian product set of action profiles that satisfies three important properties: stability against external deviations, instability against internal deviations, and minimality. This set-valued equilibrium concept generalizes the classical notion of the minimal curb set to discontinuous games. In finite games, the equilibrium cycle is related to strongly connected sink components of the best response graph.
@@
non-cooperative games,
discontinuous games,
equilibrium concept,
oscillatory game dynamics,
@@ 1 Introduction
However, there are well known issues with applying the Nash equilibrium to understand game dynamics. By game dynamics, we mean settings where players change their actions over time, seeking to improve their own payoff, in response to the actions of other players. Such game dynamics, which includes the special case of best response dynamics, do not always converge to a Nash equilibrium (Demichelis et al., 2003; Benaïm et al., 2012). From a dynamical systems standpoint, this is not surprising. After all, NEs are simply stationary points (a.k.a., equilibria) corresponding to a broad class of game dynamics, and dynamical systems do not in general converge to a stationary point. Indeed, it is not unusual for game dynamics to result in limit cycles (i.e., the action profile approaches a deterministic periodic cycle asymptotically), or even more chaotic behavior in higher dimensions; see Benaïm and Hirsch (1999); Ficici et al. (2005); Benaïm et al. (2012); Papadimitriou and Piliouras (2019). In such situations, the Nash equilibrium as a solution concept does not meaningfully capture the ‘outcome’ of the (dynamic) strategic interaction between the players.
In this paper, we introduce a novel solution concept, which we call the equilibrium cycle (EC), that seeks to capture the outcome of oscillatory game dynamics. Specifically, we restrict attention to pure (i.e., non-randomised) player strategies. The EC seeks to capture the limit set associated with a broad class of such (oscillatory) game dynamics. Crucially, the definition of the EC does not require the existence of best responses, and is therefore also applicable to discontinuous games, which arise naturally in various contexts (Dasgupta and Maskin, 1986a, b; Reny, 1999). [^1] We first discovered the equilibrium cycle while analyzing a certain pricing game between competing ride-hailing platforms (Walunj et al., 2023). This papers seeks to formalize this equilibrium notion in general terms, provide examples of equilibrium cycles in economics applications, and make connections with other classical equilibrium notions.
@@
It is easy to observe that this game does not have a pure NE (see Lotker et al. (2008)). Let us now consider natural ‘near-best response’ dynamics on this game. Note that best responses do not always exist in this game owing to discontinuities in the payoff function. By near-best response, we mean an action that yields nearly-supremal utility for a player in response to the opponent’s action.
@@ 2 Equilibrium Cycle: Definition and examples
Observe that $0<b\leq 0.5,$ since ${\mathcal{F}}$ is closed and ${\mathcal{F}}\cap{\mathcal{D}}\neq\emptyset$. We now consider the following sub-cases:
@@ 1. $(i)$
    $\{x\in[0,b)\colon(x,x)\in{\mathcal{F}}\}=\emptyset:$ Here, by the construction of $b,$ it follows that $(0,0)\notin{\mathcal{F}}$ and $(0.5,0.5)\in{\mathcal{F}},$ and thus there exists a player (w.l.o.g., say player $i$) such that $0\notin{\mathcal{F}}_{i}$ This contradicts the first condition for an EC, since the unique best response of any player $i$ to the opponent’s action $a_{-i}=0.5$ is $0\notin{\mathcal{F}}_{i}.$
@@ 2. $(ii)$
    By the construction of $b$ and $c$, it follows that $c<b,$ $(c,c)$ lies in ${\mathcal{F}},$ and $(x,x)\notin{\mathcal{F}}$ for all $x\in(c,b).$ Now, for any player $i,$ fix opponent’s action $a_{-i}=c.$ Then for $a_{i}\in(c,b),$
@@
It is instructive to note that in the classical Bertrand duopoly, $O_{c}=0,$ resulting in the unique Nash equilibrium $(c,c).$ It is the (reasonable) introduction of a positive operational cost (that is not so large that it prohibits firms from operating at all, as ensured by our assumption) that results in the EC characterized in Lemma 3. Finally, we note that the linear demand assumption is not essential to the emergence of the EC; an analogous EC can be established under any smooth decreasing demand function (so long as a positive utility is possible). The proof of Lemma 3 is in Appendix A.
The Bertrand duopoly posits an extreme ‘all or nothing’ bifurcation of market demand between the firms. The following example, which generalizes a pricing game between competing ride-hailing platforms in Walunj et al. (2023), considers a ‘less extreme’ payoff discontinuity.
@@
The game in Example 3 also admits an EC, as shown in the following lemma; this generalizes Theorem 4 in Walunj et al. (2023). The proof is once again omitted, given its similarity to the proof of Lemma 2.
@@ 5 Concluding remarks
In classical game theory, the Nash equilibrium (Nash Jr, 1950) is a foundational concept, identifying strategy profiles where no player can unilaterally deviate to improve their payoff, making it ideal to describe static one-shot interactions, and also convergent game dynamics. Indeed, when the outcome of game dynamics is a single point (strategy profile), it is typically a (pure) Nash equilibrium. However, when this is not the case, more generalized solution concepts are needed. For example, it might be meaningful to identify a set of points (strategy profiles) that become relevant in the long run, i.e., in the limit. In other words, we might anticipate that well-known game dynamics, such as best/better response dynamics, would converge to these sets over time, which can then be considered as the outcome of the game.
@@
One such set-valued outcome of game dynamics is the minimal curb set, introduced in Basu and Weibull (1991). These sets have become essential for studying adaptive processes in games, as many adjustment processes naturally settle within a minimal curb set (Voorneveld et al., 2005; Hurkens, 1995). However, this notion assumes the existence of best responses. In this paper, we define an alternative set-valued equilibrium notion, the equilibrium cycle, that does not assume the existence of best responses, and can meaningfully be applied even in discontinuous games. The equilibrium cycle is a minimal rectangular (more formally, Cartesian product) set of action profiles that is stable against external deviations, while also being unstable with respect to internal deviations. Interestingly, minimal curb sets do not always possess the latter ‘unrest’ property; however, non-trivial minimal curb sets, which exclude pure Nash equilibria, do, and are equivalent to equilibrium cycles in best response games. Importantly, payoff discontinuities arise naturally in several economics applications (Dasgupta and Maskin, 1986a, b; Reny, 1999); these discontinuous games tend not to have best responses. In this regard, the equilibrium cycle represents a significant generalization over the curb set on non-BR games, as illustrated by several examples in this paper.
@@
At the other extreme, specializing to finite games, we find that the equilibrium cycle is related to strongly connected sink components (or sink SCCs) of the best response graph. However, an exact equivalence does not hold here, owing primarily to the implicit rectangularity of equilibrium cycles (the same rectangularity does not hold in general for sink SCCs). It is possible to define a ‘non-rectangular’ variation of the equilibrium cycle that is equivalent to sink SCCs of the best response graph; however, this variant is difficult to apply to discontinuous games, and also loses the equivalence to (non-trivial, minimal) curb sets in BR games. In particular, it appears that both the rectangularity as well as the closed-ness of the EC are essential for application in discontinuous games, such as those described in Examples 13.